OpenStudy (anonymous):

Write an equation of the line that passes through (8, 3) and (4, 3). A. y = 3 B. y = 4x + 3 C. y = 4 D. y = 3x

OpenStudy (anonymous):

A!

OpenStudy (anonymous):

m=(3-3)/(8-4)=0 y-3=0(x-8) y= 3 is A